Compute the weighted average cost of capital given: The before tax cost of debt is 9.5% The cost of preferred stock is 10.75% The cost of common stock equity is 12% The tax rate is 40% The sources of capital are: Long term debt 48% Preferred stock 12% Common stock equity ????? answer using the format: 9.99%
Answer #1

Weight of Common stock equity=100-(48+12)=40%

After-tax cost of debt=9.5*(1-tax rate)

=9.5*(1-0.4)=5.7%

WACC=Respective cost*Respective weight

=(0.48*5.7)+(0.12*10.75)+(0.4*12)

=8.83%(Approx)